Solve the following initial value problem using Laplace transform. y''+16y = 2 sin 4t, y(0)=-1/2; y'(0)=0 


1
Expert's answer
2022-02-09T13:03:01-0500

"y''+16y=2\sin 4t, y(0)=-\frac{1}{2}, y'(0)=0\\"

Use Laplace transform

"L_t [y''+16y](p)=L_t[2\sin 4t](p)\\
L_t [y''](p)+L_t[16y](p)=2L_t[\sin 4t](p)\\
p^2L_t[y]-py(0)-y'(0)+16L_t[y]=2\cdot\frac{4}{p^2+16}\\
p^2L_t[y]-p(-\frac{1}{2})-0+16L_t[y]=2\cdot\frac{4}{p^2+16}\\
(p^2+16)L_t[y]=\frac{8}{p^2+16}-\frac{p}{2}\\
L_t[y]=\frac{8}{(p^2+16)^2}-\frac{p}{2(p^2+16)}\\
y=L_t^{-1}[\frac{8}{(p^2+16)^2}-\frac{p}{2(p^2+16)}](t)\\
y=L_t^{-1}[\frac{8}{(p^2+16)^2}](t)-L_t^{-1}[\frac{p}{2(p^2+16)}](t)\\
y=\frac{1}{16}(-4t\cos 4t+\sin 4t) -\frac{1}{2} \cos 4t"